OpenStudy (anonymous):

A chain store manager has been told by the main office that daily profit, P, is related to the number of clerks working that day, x, according to the equation P = -25x2 + 450x. What number of clerks will maximize the profit?

OpenStudy (anonymous):

P(x) = -25x² + 450x The graph of P(x) is a parabola opening down and has a maximum point. The extremes of a function can occur if dP(x)/dx = 0 dP(x)/dx = -50x + 450 = 0 implies the maximum occurs at x = 9.

OpenStudy (anonymous):

thanks soo very much :)